OpenStudy (anonymous):

Shoveling snow can be extremely taxing since the arms have such a low efficiency in this activity. Suppose a person shoveling a sidewalk metabolizes food at the rate of 700 W. (The efficiency of a person shoveling is 3%.) a)what is her useful power output in watts? b)how long will it take her to life 3200kg of snow in 1.20 m? c)how much waste heat will she generate in the process? (J)

OpenStudy (matt101):

When the question mentions efficiency I think it's with respect to outputing power from the metabolized food. For the first question, 700 W goes in, but only 3% of that comes out. What's the power output?

OpenStudy (anonymous):

okay so for part a i got 21 by doing 700*.03 and got it right but i don't know where to begin with the next part

OpenStudy (matt101):

Right! For the next part, remember that power is the rate of work done per unit time: P = W/t. In the second question, we want to find the t. We already have P (from the answer to the first question), so that means we need some way of finding W. Luckily, the question gives us a mass and the height it's being lifted. You have to do work to lift that mass...and the work done is equal to the potential energy gained by that mass (given by mgh). Does that help you out?

OpenStudy (anonymous):

YES! So i did 3200*9.8*1.2 and got 37632, then i rearranged p=w/t to t=w/p and got 1792 seconds (which is correct). How do i determine waste heat?

OpenStudy (matt101):

Well the waste heat is all the energy that went in that wasn't converted into useful work. So the waste heat "power" is 700-21. Use the this and the time you just calculated to calculate the "work", which is really the energy lost as a result of heat.

OpenStudy (anonymous):

Thank you so much you are a life saver!

OpenStudy (matt101):

No problem!